OpenStudy (anonymous):

PLEASE HELP! 13 = 5 + 3b - 13

OpenStudy (anonymous):

b=-5/3

OpenStudy (anonymous):

How'd U get that? please I need the work shown. Thank you

OpenStudy (anonymous):

26 = 5 + 3b 21 = 3b 7 = b

OpenStudy (anonymous):

THANKS!!! (: